# Private Fine-tuning of Large Language Models with Zeroth-order Optimization

### v2

@@ Abstract
Differentially private stochastic gradient descent (DP-SGD) allows models to be trained in a privacy-preserving manner, but has proven difficult to scale to the era of foundation models. We introduce DP-ZO, a private fine-tuning framework for large language models by privatizing zeroth order optimization methods. A key insight into the design of our method is that the direction of the gradient in the zeroth-order optimization we use is random and the only information from training data is the step size, i.e., a scalar. Therefore, we only need to privatize the scalar step size, which is memory-efficient. DP-ZO provides a strong privacy-utility trade-off across different tasks, and model sizes that are comparable to DP-SGD in $(\varepsilon,\delta)$-DP. Notably, DP-ZO possesses significant advantages over DP-SGD in memory efficiency, and obtains higher utility in $\varepsilon$-DP when using the Laplace mechanism.
@@ 1 INTRODUCTION
Figure 1: Visualization of DP-ZO. The only information from private data is a scalar and we only need to add noise to this scalar. This scalar privatization enjoys the benefits of flexibility with DP mechanisms, ease of implementation, and reduced computation.
The proliferation of open-source models pretrained on web-scale datasets (Brown et al., 2020; Zhang et al., 2022; Touvron et al., 2023) has created a paradigm shift in privacy preserving machine learning. Differential Privacy (DP) (Dwork et al., 2006) is the gold standard for preserving privacy while training models on private data, but it requires additional data (Tramèr and Boneh, 2021) to prevent a drop in utility (Yu et al., 2021a). Pretrained model checkpoints have emerged as a compelling “free” source of prior information to boost the performance of DP training (Ganesh et al., 2023; Tang et al., 2023a; Panda et al., 2024a). By only requiring DP during the fine-tuning phase, a recent line of work (Li et al., 2022b, a; Yu et al., 2022; He et al., 2023; Bu et al., 2023d) is able to obtain impressive performance with DP-SGD (Abadi et al., 2016). Despite these advancements, DP-SGD causes additional memory cost and needs additional engineering effort, especially for large models across devices. We propose a new direction for DP fine-tuning of large pretrained models that achieves strong privacy-utility trade-off and is more resource-efficient, easy to implement, and portable.
In this work, we introduce a new methodology DP-ZO for DP fine-tuning of large pretrained models. Our method uses zeroth-order optimization (ZO) (Spall, 1992). Our key insight is the synergy between differentially private fine-tuning and zeroth-order optimization. ZO provides the gradient estimates and the only information from private data in ZO is a scalar. We only need to privatize the scalar update by adding noise to it. Specifically, the scalar is the differences between losses from models with the same random perturbation but flipped signs. DP-ZO privatizes the zeroth-order update, by adding noise to the difference between the losses (visualized in Figure 1). This noise is proportional to the sensitivity of this loss difference with respect to changing a single example in the training set, which is controlled by clipping. We limit the $\ell_{p}$ sensitivity by clipping the norm of the difference in scalar losses, between the two random perturbations. Therefore, DP-ZO is flexible for both $\varepsilon$-DP and $(\varepsilon,\delta)$-DP. By removing the need for per-example gradient clipping (Abadi et al., 2016), DP-ZO enables DP training of language models with just a few lines of code and without the need for backpropagation.
Figure 2: DP-ZO provides a strong privacy-utility trade-off across different tasks under conservative privacy budgets.
@@
Besides, we provide the empirical privacy analysis of ZO and DP-ZO. While ZO itself incurs less empirical privacy loss than SGD, such empirical privacy analysis is still much higher than random guess. DP-ZO can reduce such privacy attack close to random guess. We also show the computation efficiency of DP-ZO over DP-SGD, even when applying gradient checkpointing and half-precision to both methods.
@@ 2.1 Differential Privacy
An algorithm $\mathcal{M}$ is said to be $(\varepsilon,\delta)$-DP if for all sets of events $S\subseteq\text{Range}(\mathcal{M})$ and neighboring datasets $D,D^{\prime}\in\mathcal{D}^{n}$ (where $\mathcal{D}$ is the set of all possible data points) we have the guarantee:
@@
We define a set of existing DP mechanisms that we will use in our work.
Proposition 2.2 (Gaussian mechanism (Dwork and Roth, 2014)).
@@
where $z\sim\mathcal{N}\left(0,\frac{2\ln(1.25/\delta)\Delta^{2}}{\varepsilon^{2}}\right)$, provides $(\varepsilon,\delta)$-DP.
Proposition 2.3 (Laplace mechanism (Dwork and Roth, 2014)).
@@ 2.2 Zeroth-order Optimization
We use a method from the large body of work on zeroth-order optimization (Flaxman et al., 2004; Shamir, 2013; Ghadimi and Lan, 2013; Nesterov and Spokoiny, 2017) that uses the difference in losses between two random perturbations (Duchi et al., 2015; Spall, 1992) with opposite signs to determine the magnitude of a gradient update in the direction of the random perturbations. In the non-private setting where the adaptation between the pretrained model and the fine-tuning dataset has low rank (Hu et al., 2022), as in fine-tuning large language models, Malladi et al. (2023) show this method converges at a rate that is not catastrophically slower than SGD fine-tuning.
Definition 2.4 (Simultaneous Perturbation Stochastic Approximation (SPSA)  (Spall, 1992))).
Given a model with parameters $\theta\in\mathbb{R}^{d}$ and a loss function $\mathcal{L}$, the gradient estimate on a minibatch drawn from a dataset $\mathcal{B}\subset\mathcal{D}$ is computed by projecting the loss on the minibatch $\mathcal{L}(\theta;\mathcal{B})$ onto a random perturbation $z\in\mathbb{R}^{d}$ that is a standard Gaussian random variable (i.e., $z\sim\mathcal{N}(0,I_{d})$) scaled by $\phi$:
@@
Random perturbations in zeroth-order optimization (ZO) serve as high-variance estimates of the actual gradient, enabling optimization without the need for explicit gradient computations. However, these perturbations themselves carry a privacy risk. The characteristics of the perturbations can be inferred from the gradient updates, effectively leaking information about the data. Therefore, incorporating differential privacy into ZO is essential to safeguard against these vulnerabilities.
@@ 3 OUR METHOD: DP-ZO
We introduce our framework for differentially private zeroth order optimization (DP-ZO) by integrating DP into Definition 2.4. In our framework, the information obtained from training data can be represented as a scalar. This scalar has a bounded sensitivity (when applying clipping) and can be privatized by adding noise. If we compare the noise added in DP-ZO to a single dimension to the noise added in DP-SGD to the entire gradient, we expect the univariate noise to be less detrimental to the utility (due to the curse dimensionality in differential privacy Dwork and Roth (2014)). In other words, we would expect the gap between non-private and private utility to be smaller than that of DP-SGD. However, it is possible that for some tasks the non-private performance of zeroth-order optimization is poor (see Section 4.3).
@@
1:Model parameters $\theta$, dataset $\mathcal{D}$, learning rate $\eta$, perturbation scale $\phi$, privacy parameter $\sigma$, noising mechanism $\mathcal{Z}$, clipping threshold ${C}$, expected batch size $B$, sub-sampling rate $p=B/|\mathcal{D}|$).
2:$g={0}$
3:for $t\in 1,\ldots T$ do
4:     Poisson sample $\mathcal{B}$ from $\mathcal{D}$ with sub-sampling rate $p$
5:     $\vec{z}\sim\mathcal{N}(\vec{0}_{|\theta|},\mathbf{I}_{|\theta|\times|\theta|})$
6:     $\theta^{+}\leftarrow\theta+\phi\vec{z}$
7:     $\theta^{-}\leftarrow\theta-\phi\vec{z}$
8:     for $(x_{i},y_{i})\in\mathcal{B}$ do
9:         $l_{i}^{+}\leftarrow\mathcal{L}(\theta^{+},(x_{i},y_{i}))$
10:         $l_{i}^{-}\leftarrow\mathcal{L}(\theta^{-},(x_{i},y_{i}))$
11:         $l_{i}=clip(l_{i}^{+}-l_{i}^{-},C)$
12:     end for
13:     $s=\frac{\sum_{i\in\mathcal{B}}l_{i}+\mathcal{Z}(\mathrm{C},\sigma)}{B\cdot 2\phi}$
14:     $\theta=\theta-\eta s\vec{z}$
15:end for
@@
DP-ZO. We explain the steps of our algorithm while emphasizing the key differences from Definition 2.4 required to guarantee $(\varepsilon,\delta)$-DP. We first sample a batch from the dataset with Poisson sampling (Balle et al., 2018) which allows us to use privacy amplification by subsampling. For each model parameter $\theta_{i}$ we want to update, we independently sample a perturbation $z_{i}$ from a standard Gaussian distribution and scale it by a predetermined constant $\phi$; we denote the full perturbation vector as $\phi\vec{z}$. Now we compute an approximation of the gradient by projecting it onto the random perturbation $\vec{z}$. That is, for a training sample $x_{i}$ we compute the difference in scalar losses between ${\theta+\phi\vec{z},\theta-\phi\vec{z}}$. Intuitively, this scalar tells us how much better one random step is than the other. We clip this scalar to limit the sensitivity . We add noise to the aggregation over samples in our training batch (described in detail in the subsequent paragraph). Finally, we take a step in the direction of $\vec{z}$ by scaling our private step size by the expected batch size, perturbation constant $\phi$, and the learning rate $\eta$.
@@
DP-ZO enables new mechanisms by privatizing the difference in losses between perturbations. DP-ZO proposes an update direction determined by a $d$-dimensional random vector (sampled from standard Gaussian distribution) independent of the private training data. The only private aspect is the step size, that is influenced by the difference in losses between perturbations with opposite signs. To privatize this step size, we add noise proportional to the sensitivity of the step size. We bound the sensitivity of the step size by clipping the per-example step sizes to a specific range $[-C,C]$, so the sensitivity under add-remove DP is $C$.
@@
Given a private scalar with bounded sensitivity, we can apply the classical Gaussian mechanism to release a privatized scalar with $(\varepsilon,\delta)$-DP. The Gaussian mechanism is widely studied in privacy-preserving machine learning techniques like DP-SGD, in part because the best accounting techniques for the Gaussian mechanism (Dong et al., 2019; Gopi et al., 2021) are tight. However, the Gaussian mechanism can only provide $(\varepsilon,\delta)$-DP and researchers often recommend using cryptographically small values of $\delta$ (Vadhan, 2017). Unfortunately, due to limitations of accounting methods, we currently cannot calculate the tight privacy of composition of sub-sampled Gaussian mechanism for values of $\delta$ smaller than $10^{-10}$. Alternatively, we can resort to mechanisms that can obtain pure $\varepsilon$-DP. These mechanisms, such as Laplace mechanism, come with a guarantee that the mechanism will never fail catastrophically. However, due to large tails of the Laplace mechanism, it has never been a contender for high dimensional optimization.
@@
Although it is possible to obtain pure DP with DP-SGD by adding Laplace noise scaled to the $\ell_{1}$ sensitivity of the gradient, this is challenging for large models because the $\ell_{1}$ sensitivity can be $\sqrt{d}$ times larger than the $\ell_{2}$ sensitivity (and often is; see Section 4.3), especially for billion-parameter LLMs. In contrast, DP-ZO only requires privatizing the loss. The one-dimensional private estimation of the step size is amenable to the Laplace mechanism, because the $\ell_{p}$ norms are equivalent. Specifically, DP-ZO with the Laplace mechanism is the first method to achieve a reasonable privacy-utility trade-off under pure $\epsilon$-DP for private fine-tuning of LLMs. While this work primarily explores these two mechanisms, the DP-ZO framework is flexible enough to be extended to other differential privacy mechanisms, broadening its applicability.
@@
Algorithm 1 is $(\varepsilon,\delta)$-DP.
Proof Overview. We analyze the privacy of models trained iteratively with DP-ZO by composing the per-iteration privacy loss (Kairouz et al., 2015). At each step, line 11 upper bounds the $\ell_{1}$ (and therefore $\ell_{p}\forall p\geq 1$) sensitivity of the difference in losses $l_{i}$ by $C$. Line 13 adds noise based on DP mechanisms such that each step in Algorithm 1 satisfies $(\varepsilon,\delta)$-DP with some privacy parameters. Therefore Algorithm 1 is $(\varepsilon,\delta)$-DP with some privacy parameters that are calculated via some mechanism-dependent composition theorem. The specific $\varepsilon,\delta$ depends on the choice of mechanism. Most of the privacy analyses in Appendix A are based on the numerical composition of private random variable for the corresponding privacy curve (Gopi et al., 2021) except for the pure $\varepsilon$-DP analysis. The analysis for pure $\varepsilon$-DP by the Laplace mechanism is based on the basic composition (Dwork and Roth, 2014). Given a number of iterations $T$ and the use of the Laplace mechanism $Laplace(0,\sigma)$, Algorithm 1 is $\varepsilon$-DP with $\varepsilon=T\cdot\log(1+p\cdot(e^{1/\sigma}-1))$.
@@
Proposition 3.2.
DP-ZO attains a convergence rate $\mathcal{O}(\sqrt{r}/\varepsilon n)$, where $r$ is the effective rank of the problem.
Malladi et al. (2023) proves the convergence rate of fine-tuning of language models with zeroth-order optimization is proportional to $r$ instead of the model dimension. A concurrent work (Zhang et al., 2024a) independently proposes DP-ZO and provides the convergence analysis for DP-ZO with Gaussian mechanism, that is independent of the model dimension in private training (Song et al., 2021; Li et al., 2022a). We discuss our work and Zhang et al. (2024a) in Section 6.

### v3

@@ Abstract
Differentially private stochastic gradient descent (DP-SGD) allows models to be trained in a privacy-preserving manner, but has proven difficult to scale to the era of foundation models. We introduce DP-ZO, a private fine-tuning method for large language models by privatizing zeroth order optimization methods. A key insight into the design of our method is that the direction of the gradient in the zeroth-order optimization we use is random and the only information from the training data is the step size, i.e., a scalar. Therefore, we only need to privatize the scalar step size, which is memory-efficient. DP-ZO provides a strong privacy-utility trade-off across different tasks, and model sizes that are comparable to DP-SGD in $(\varepsilon,\delta)$-DP. Notably, DP-ZO possesses significant advantages over DP-SGD in memory efficiency, and obtains higher utility in pure $\varepsilon$-DP when using the Laplace mechanism.
@@ 1 INTRODUCTION
Figure 1: Visualization of DP-ZO. The only information from private data is a scalar step size for direction with lower target function value and we only need to add noise to this scalar. This scalar privatization enjoys the benefits of flexibility with DP mechanisms, ease of implementation, and reduced computation.
The proliferation of open-source models pretrained on web-scale datasets (Brown et al., 2020; Zhang et al., 2022; Touvron et al., 2023) has created a paradigm shift in privacy preserving machine learning. Differential Privacy (DP) (Dwork et al., 2006) is the gold standard for preserving privacy while training models on private data, but it requires additional data (Tramèr & Boneh, 2021) to prevent a drop in utility (Yu et al., 2021a). Pretrained model checkpoints have emerged as a compelling “free” source of prior information to boost the performance of DP training (Ganesh et al., 2023; Tang et al., 2023; Panda et al., 2024a). By only requiring DP during the fine-tuning phase, a recent line of work (Li et al., 2022b; a; Yu et al., 2022; He et al., 2023; Bu et al., 2023d) is able to obtain impressive performance with DP-SGD (Abadi et al., 2016). Despite these advancements, DP-SGD causes additional memory cost and needs additional engineering effort, especially for large models across devices. We propose a new direction for DP fine-tuning of large pretrained models that achieves strong privacy-utility trade-off and is more resource-efficient, easy to implement, and portable.
In this work, we study DP fine-tuning of large pretrained models with zeroth-order optimization and introduce DP-ZO. Our method uses zeroth-order optimization (ZO) (Spall, 1992). Our key insight is the synergy between differentially private fine-tuning and zeroth-order optimization. ZO provides the gradient estimates and the only information from private data in ZO is a scalar. We only need to privatize the scalar update by adding noise to it. Specifically, the scalar is the differences between losses from models with the same random perturbation but flipped signs. DP-ZO privatizes the zeroth-order update, by adding noise to the difference between the losses (visualized in Figure 1). This noise is proportional to the sensitivity of this loss difference with respect to changing a single example in the training set, which is controlled by clipping. We limit the $\ell_{p}$ sensitivity by clipping the norm of the difference in scalar losses, between the two random perturbations. Therefore, DP-ZO is flexible for both $\varepsilon$-DP and $(\varepsilon,\delta)$-DP. By removing the need for per-example gradient clipping (Abadi et al., 2016), DP-ZO enables DP training of language models with just a few lines of code without backpropagation.
Figure 2: DP-ZO provides a strong privacy-utility trade-off across different tasks under conservative privacy budgets. F1 is for SQuAD and DROP and accuracy is for SST2.
@@
Besides, we provide the empirical privacy analysis of (DP-)ZO. While ZO itself incurs less empirical privacy loss than SGD, such empirical privacy leakage estimated by membership inference attacks (Shokri et al., 2017; Panda et al., 2024b) is still much higher than random guess. Compared to ZO, DP-ZO can reduce this membership inference attack close to random guess. We also show the computation efficiency of DP-ZO over DP-SGD, even when applying gradient checkpointing and half-precision to both methods.
Independently and concurrently, Zhang et al. (2024a) also studied privatizing the scalar loss in Zeroth-order optimization with Gaussian noise and proposed DPZero. Our DP-ZO shares some similar motivation and design as DPZero, and there are several differences between our work and theirs. We provide a detailed discussion on DP-ZO and Zhang et al. (2024a) in Section 5.
@@ 2.1 Differential Privacy
We call $D,D^{\prime}$ are neighboring datasets if they differ in exactly one record by adding or removing one record. [^1] This neighboring definition is for add/removal DP and the most common one. There are other neighboring definitions in the literature (Desfontaines & Pejó, 2020; Ponomareva et al., 2023). An algorithm $\mathcal{M}$ is said to be $(\varepsilon,\delta)$-DP if for all sets of events $S\subseteq\text{Range}(\mathcal{M})$ and neighboring datasets $D\simeq D^{\prime}$ we have the guarantee:
@@
For a vector $x=(x_{1},x_{2},..,x_{n})$, we use $l_{p}$ to denote the p-norm of $x$, $l_{p}(x)=(|x_{1}|^{p}+|x_{2}|^{p}+...+|x_{n}|^{p})^{\frac{1}{p}}$. We now introduce a set of existing DP mechanisms that we will use in our work.
Proposition 2.2 (Gaussian mechanism (Dwork & Roth, 2014; Balle & Wang, 2018)).
@@
where $z\sim\mathcal{N}\left(0,\sigma^{2}\right)$, provides $(\varepsilon,\delta)$-DP where $\Phi(\frac{\Delta}{2\sigma}-\frac{\varepsilon\sigma}{\Delta})-e^{\varepsilon}\Phi(-\frac{\Delta}{2\sigma}-\frac{\varepsilon\sigma}{\Delta}))\leq\delta$. $\Phi(t)$ is the cumulative distribution function (CDF) of the univariate Gaussian distribution $\mathcal{N}(0,1)$.
Proposition 2.3 (Laplace mechanism (Dwork & Roth, 2014)).
@@ 2.2 Zeroth-order Optimization
Zeroth-order optimization methods (Kiefer & Wolfowitz, 1952; Spall, 1992; Shamir, 2013; Ghadimi & Lan, 2013; Nesterov & Spokoiny, 2017) use finite difference of function values to estimate gradients, instead of computing gradients in first-order methods like SGD. By evaluating the objective function values around the points $x$, ZO provides the step size towards the direction where the point has a lower function value; See Liu et al. (2020); Zhang et al. (2024c) for a more detailed review of Zeroth-order optimization methods. Particularly, we use the difference in losses between two random perturbations SPSA (Spall, 1992; Duchi et al., 2015) with opposite signs to determine the magnitude of a gradient update in the direction of the random perturbations. Following (Spall, 1992; Malladi et al., 2023), we define SPSA in Definition 2.4.
Definition 2.4 (Simultaneous Perturbation Stochastic Approximation (SPSA) (Spall, 1992)).
Given a model with parameters $\theta\in\mathbb{R}^{d}$ and a loss function $\mathcal{L}$, the gradient estimate on a minibatch $\mathcal{B}$ drawn from a dataset $\mathcal{D}$ is computed by projecting the loss on the minibatch $\mathcal{L}(\theta;\mathcal{B})$ onto a random perturbation $z\in\mathbb{R}^{d}$ that is a standard Gaussian random variable (i.e., $z\sim\mathcal{N}(0,I_{d})$) scaled by perturbation scale $\phi$:
@@
As noted in the Malladi et al. (2023), when $\phi\to 0$, the SPSA estimate could be considered as a rank-1 reconstruction of the gradient. While SPSA only provides a scalar information from the data, interestingly,  Malladi et al. (2023) show this method converges at a rate that is not catastrophically slower than SGD in fine-tuning large language models in downstream tasks. Malladi et al. (2023) reason this phenomena as a result of the Hessian of the loss exhibiting small local effective rank.
Zeroth-order optimization (ZO) serves as high-variance estimates of the actual gradient (Liu et al., 2018), enabling optimization without the need for explicit gradient computations. While the update of model is from the random perturbation scaled by the step size and the only information from data is the step size, ZO still carries a privacy risk, leaking information about the data (as we show later in Section 4.3). Therefore, incorporating differential privacy into ZO is essential to safeguard against these vulnerabilities.
@@ 3 OUR METHOD: DP-ZO
We introduce our method for differentially private zeroth order optimization (DP-ZO) by integrating DP into Definition 2.4. In DP-ZO, the information obtained from training data can be represented as a scalar. This scalar has a bounded sensitivity (when applying clipping) and can be privatized by adding noise. If we compare the noise added in DP-ZO to a single dimension to the noise added in DP-SGD to the entire gradient, we expect the univariate noise to be less detrimental to the utility (due to the curse dimensionality in differential privacy (Dwork & Roth, 2014)). In other words, we would expect the gap between non-private and private utility to be smaller than that of DP-SGD. However, it is possible that for some tasks the non-private performance of zeroth-order optimization is poor (see Section 4.3).
@@
1:Model parameters $\theta$, dataset $D$, learning rate $\eta$, perturbation scale $\phi$, privacy parameter $\sigma$, noising mechanism $\mathcal{Z}$ (Gaussian or Laplace), clipping threshold ${C}$, expected batch size $B$, sub-sampling rate $p=B/|D|$).
2:for $t\in 1,\ldots T$ do
3:     Poisson sample $\mathcal{B}$ from $D$ with sub-sampling rate $p$
4:     $\vec{z}\sim\mathcal{N}(\vec{0}_{|\theta|},\mathbf{I}_{|\theta|\times|\theta|})$
5:     $\theta^{+}\leftarrow\theta+\phi\vec{z}$
6:     $\theta^{-}\leftarrow\theta-\phi\vec{z}$
7:     for $(x_{i},y_{i})\in\mathcal{B}$ do
8:         $l_{i}^{+}\leftarrow\mathcal{L}(\theta^{+},(x_{i},y_{i}))$
9:         $l_{i}^{-}\leftarrow\mathcal{L}(\theta^{-},(x_{i},y_{i}))$
10:         $l_{i}=clip(l_{i}^{+}-l_{i}^{-},C)$
11:     end for
12:     if $\mathcal{Z}$ is Gaussian then
13:         $s=\frac{\sum_{i\in\mathcal{B}}l_{i}+\mathcal{N}(0,C^{2}\sigma^{2})}{B\cdot 2\phi}$
14:     else if $\mathcal{Z}$ is Laplace then
15:         $s=\frac{\sum_{i\in\mathcal{B}}l_{i}+Laplace(0,C\sigma)}{B\cdot 2\phi}$
16:     end if
17:     $\theta=\theta-\eta s\vec{z}$
18:end for
@@
DP-ZO. We explain the steps of our algorithm while emphasizing the key differences from Definition 2.4 required to provide DP guarantee. We first sample a batch from the dataset with Poisson sampling (Balle et al., 2018) which allows us to use privacy amplification by subsampling. For each model parameter $\theta_{i}$ we want to update, we independently sample a perturbation $z_{i}$ from a standard Gaussian distribution and scale it by a predetermined constant $\phi$; we denote the full perturbation vector as $\phi\vec{z}$. Now we compute an approximation of the gradient by projecting it onto the random perturbation $\vec{z}$. That is, for a training sample $x_{i}$ we compute the difference in scalar losses between ${\theta+\phi\vec{z},\theta-\phi\vec{z}}$. Intuitively, this scalar tells us how much better one random step is than the other. We clip this scalar to limit the sensitivity . We add noise to the aggregation over samples in our training batch (described in detail in the subsequent paragraph). Finally, we take a step in the direction of $\vec{z}$ by scaling our private step size by the expected batch size, perturbation constant $\phi$, and the learning rate $\eta$.
@@
DP-ZO enables new mechanisms by privatizing the difference in losses between perturbations. DP-ZO proposes an update direction determined by a $d$-dimensional random vector (sampled from standard Gaussian distribution) independent of the private training data. The only private information is the step size, that is influenced by the difference in losses between perturbations with opposite signs. To privatize this step size, we add noise proportional to the sensitivity of the step size. We bound the sensitivity of the step size by clipping the per-example step sizes to a specific range $[-C,C]$, so the sensitivity under add-remove DP is $C$.
@@
Given a private scalar with bounded sensitivity, we can apply the classical Gaussian mechanism to release a privatized scalar with $(\varepsilon,\delta)$-DP. The Gaussian mechanism is widely studied in privacy-preserving machine learning techniques like DP-SGD. However, the Gaussian mechanism can only provide $(\varepsilon,\delta)$-DP and researchers often recommend using cryptographically small values of $\delta$ (Vadhan, 2017). The stronger privacy notion, i.e., $\varepsilon$-DP, comes with a guarantee that the mechanism will never fail catastrophically. For example, we can use the Laplace mechanism to achieve the pure $\varepsilon$-DP guarantee. However, due to large tails of the Laplace mechanism, it has never been a contender for high dimensional optimization.
@@
Although it is possible to obtain pure DP with DP-SGD by adding Laplace noise scaled to the $\ell_{1}$ sensitivity of the gradient, this is challenging for large models because the $\ell_{1}$ sensitivity can be $\sqrt{d}$ times larger than the $\ell_{2}$ sensitivity (and often is; see Section 4.3), especially for billion-parameter LLMs. In contrast, DP-ZO only requires privatizing the loss. The one-dimensional private estimation of the step size is amenable to the Laplace mechanism, because the $\ell_{p}$ norms are equivalent for the scalar. Specifically, DP-ZO with the Laplace mechanism is the first method to achieve a reasonable privacy-utility trade-off under pure $\epsilon$-DP for private fine-tuning of LLMs. While this work primarily explores these two mechanisms, DP-ZO is flexible enough to be extended to other differential privacy mechanisms, broadening its applicability.
@@
Algorithm 1 is $(\varepsilon,\delta)$-DP. Particularly, for Laplace mechanism $Laplace(0,\sigma)$, Algorithm 1 is $\varepsilon$-DP with $\varepsilon=T\cdot\log(1+p\cdot(e^{1/\sigma}-1))$.
Proof Overview. We first provide the privacy analysis for each step. At each step, line 10 upper bounds the $\ell_{1}$ (and therefore $\ell_{p}\forall p\geq 1$ for the scalar value) sensitivity of the difference in losses $l_{i}$ by $C$. Lines 13 and 15 add noise based on DP mechanisms such that each step in Algorithm 1 satisfies $(\varepsilon,\delta)$-DP with some privacy parameters. For Laplace mechanism, each step satisfies $\varepsilon$-DP where $\varepsilon=\log(1+p\cdot(e^{1/\sigma}-1))$. For Gaussian mechanism, each step satisfies $(\varepsilon,\delta)$-DP where $\text{max}(H_{e^{\varepsilon}}(\mathcal{N}(0,\sigma^{2})\|(1-p)\mathcal{N}(0,\sigma^{2})+p\mathcal{N}(1,\sigma^{2})),H_{e^{\varepsilon}}((1-p)\mathcal{N}(1,\sigma^{2})+p\mathcal{N}(0,\sigma^{2})\|\mathcal{N}(1,\sigma^{2})))\leq\delta$ and $H_{e^{\varepsilon}}(\cdot\|\cdot)$ is Hockey-stick divergence (see Definition A.3 in Appendix A). Therefore Algorithm 1 is $(\varepsilon,\delta)$-DP with some privacy parameters that are calculated via some mechanism-dependent composition theorem. Most of the privacy analyses in Appendix A are similar to the privacy analysis of DP-SGD and are based on the numerical composition of privacy loss random variable (see Definition A.7 in Appendix A) for the corresponding privacy curve (Gopi et al., 2021) except for the pure $\varepsilon$-DP analysis. The analysis for pure $\varepsilon$-DP by the Laplace mechanism is based on the basic composition (Dwork & Roth, 2014). Given a number of iterations $T$ and the use of the Laplace mechanism $Laplace(0,C\sigma)$, Algorithm 1 is $\varepsilon$-DP with $\varepsilon=T\cdot\log(1+p\cdot(e^{1/\sigma}-1))$.
@@
Prior works (Song et al., 2021; Li et al., 2022a) have analyzed when the convergence rate of DP-SGD is dependent on the effective rank $r$ of the problem rather than the model dimension size $d$. A concurrent and independent work (Zhang et al., 2024a) propose DPZero by privatizing the loss difference scalar in ZO with Gaussian noise, that shares some similar design as our DP-ZO. They provide the convergence guarantee for DPZero that is independent of the model dimension in private training. Here we briefly discuss the algorithm design difference between DPZero in Zhang et al. (2024a) and our DP-ZO. DPZero is Algorithm 2 in Zhang et al. (2024a), and the random perturbation is by sampling $\vec{z}$ from a unit Sphere $\mathbb{S}^{d-1}=\{x\in\mathbb{R}^{d}|\lVert x\rVert=1\}$. In our Algorithm 1, $\vec{z}$ is sampled from Gaussian distribution. We reimplemented our perturbation method based on the algorithms in Zhang et al. (2024a), and we obtain $82.32_{0.82}$ for LoRA fine-tuning OPT-13B on SQuAD with $(1,10^{-5})$-DP, that is comparable as the perturbation method in our Algorithm 1, that achieves $82.28_{0.84}$ under the same setting. We also note that Zhang et al. (2024a) also share same observation on this by experiments on Roberta-large models (Liu et al., 2019) in their Table 7. Algorithm 2 in Zhang et al. (2024a) for DPZero is under full batch setting. [^2] The experiment section in Zhang et al. (2024a) also uses small batch size and perturbation from random Gaussian distribution. As we show in Section D.2, in the evaluated number of iteration range, given the same amount of computation time on a single GPU, by training more steps with small batch size, DP-ZO achieves better performance than by accumulating steps for large batch size. This result is consistent with the observation in the non-private zeroth-order optimization results in MeZO (Malladi et al., 2023). Besides, Algorithm 2 in DPZero considers the sensitivity as $2C$ with clipping threshold $C$, this leads to adding twice necessary noise under add/remove DP. Our Algorithm 1 analyzes the sensitivity $C$ with clipping threshold $C$, and we share this add/remove DP choice and sensitivity $C$ with popular DP-SGD library like Opacus (Yousefpour et al., 2021). We further discuss our work and Zhang et al. (2024a) in Section 5.